OpenStudy (anonymous):

Solve for x: -2/3x+1/5x=2 A. -1/ 10 B. -7/ 30 C. -4/ 15 D. 13/ 30 E. None of the above

OpenStudy (anonymous):

Im so confused because the answer should be -7/30 or -30/7

OpenStudy (anonymous):

I get -30/7

OpenStudy (anonymous):

did you get that by using the reciprocal?

OpenStudy (anonymous):

\[-\frac{ 2 }{ 3 }x+\frac{ 1 }{ 5 }x=2\] \[-\frac{ 10 }{ 15}x+\frac{ 3 }{ 15}x=2\] becomes \[-\frac{ 7 }{ 15 }x=2\] \[x=2*-\frac{ 15 }{ 7 }\] \[x=-\frac{ 30 }{ 7 }\]

OpenStudy (anonymous):

if you plug that in for x it does not equal 2.-7/30 will equal 2

OpenStudy (anonymous):

the x is in the denominator

OpenStudy (anonymous):

I did the math and got 2

OpenStudy (anonymous):

are you using an calculator? if so make sure you use ()

OpenStudy (anonymous):

a*

OpenStudy (anonymous):

n. How would -30/7 if you have negative 7 over 15x equals two.

OpenStudy (anonymous):

I don't have -7 over 15. I have -15 over 7

OpenStudy (anonymous):

in this problem the x is in the denominator. And I am not sure how to isolate x

OpenStudy (anonymous):

Are you sure the x is in the denom and not being multiplied by the whole fraction?

OpenStudy (anonymous):

can you use the equation tab below the chat box to type in the equation exactly the way it look in the book.

OpenStudy (anonymous):

yes and that is what has me stuck. is proper \[\frac{ -3 }{3x }+\frac{ 1 }{ 5x }=2\]

OpenStudy (anonymous):

oh okay.

OpenStudy (anonymous):

-2/3x

OpenStudy (anonymous):

is it -3/3x or -2/3x?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[\frac{ -2 }{ 3x }+\frac{ 1 }{ 5x }=2\]

OpenStudy (anonymous):

I am sorry. I'm stumped. bump the question and perhaps someone else can help you.

OpenStudy (anonymous):

I do not see a way to get -7/30 out of the original equation.

OpenStudy (anonymous):

don't you need to multiply 15x by the two also to keep both sides equal

OpenStudy (anonymous):

you get -10 +3 = 30x then divide by 30

OpenStudy (anonymous):

yes when you multiply 2 by 15x you get 30x then you divide the 30

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

okay, im glad you got it. np